# Rehearsal-free Federated Domain-incremental Learning

### v1

@@ Abstract
1 Introduction
@@
Federated Learning (FL) is a remarkable approach that enables various participants, also referred to as clients, to collaboratively train a global machine learning model while keeping their data decentralized, maintaining privacy and leveraging diverse data sources [24, 9]. A significant limitation of current FL research lies in its primary focus on scenarios with static data distributions [20, 12, 17, 13, 34], which remain unchanged throughout the training process [8, 7]. This limitation is particularly pronounced as it does not reflect the dynamic nature of real-world data. Recent studies have therefore shifted towards Federated Continual Learning (FCL), especially in class-incremental data scenarios [8, 7, 38, 23, 2], which struggles significantly with the critical problem of catastrophic forgetting [30], where model rapidly forgets previously learned knowledge upon encountering new data. In this paper, we explore more challenging domain-incremental data scenarios.
Figure 1: Key steps of RefFil framework. Panel (a) shows the common setting in existing FCL works, characterized by a cliff-style data transition. Panel (b) shows our approach, where a subset of participants gradually transitions to new tasks. Panel (c) provides an overview example of a key step in our methodology: the 1st participant processes new domain data using global prompts from the 2nd to $m$-th participants and local prompts, enhancing robustness by aligning the model’s predictions across diverse domain prompts as inputs.
@@
Various methods have been proposed to address catastrophic forgetting. Rehearsal-based approaches, such as those described in [30], mitigate forgetting by replaying a selection of old task data during new task training. Alternatively, regularization-based methods, like Knowledge Distillation (KD) [21, 30, 3, 8], guide the model’s optimization using historical data or models. Network expansion techniques [18, 11, 29] adapt the model structure to incorporate new features but their practicality in memory-constrained FL environments is limited. Recent research like L2P [37] and DualPrompt [36] explores prompt-based, rehearsal-free domain-incremental learning. However, these methods still do not fully address the challenge of using global domain information to enhance local model robustness in Federated Domain-Incremental Learning (FDIL).
@@
Strategies for mitigating catastrophic forgetting in domain-incremental tasks are distinct from those employed in class-incremental scenarios, where the substantial differences between classes make it easier for models to distinguish decision boundaries. In domain-incremental contexts, however, the smaller feature differences between domains make it harder to distinguish decision boundaries. Thus, the priority shifts towards acquiring domain-invariant knowledge, an effort pivotal to enhancing model robustness. This focus becomes particularly vital when models are required to make predictions across a spectrum of diverse domains. To date, research in this area is notably sparse, especially concerning the effective global utilization of heterogeneous domain information from all clients at various stages of their transition. Such utilization is key to mitigating catastrophic forgetting by strengthening the local models’ proficiency in learning domain-invariant knowledge in FDIL and across the broader expanse of FCL.
@@
2 Related Work
@@
Most recently, several works have begun to concentrate on FCL, especially Federated Class-incremental Learning (FCIL), which involves sequentially learning a series of distinct, non-overlapping classes. The primary challenge in FCIL is catastrophic forgetting, where models tend to lose previously learned knowledge upon task transition. Approaches such as GFLC [8] and its extension LGA [7] use knowledge distillation, employing an old model to regularize the current model’s output and a loss function to mitigate global forgetting. Method TARGET [38] uses a global model to generate synthetic data with a global distribution for the current training task to reduce forgetting. CFed [23] addresses data unavailability by generating pseudo-labels on a surrogate dataset for knowledge distillation and proposes a server distillation mechanism to mitigate intra-task forgetting. These methods typically require storing old tasks or extra data, which is impractical in FL with resource-limited devices. MFCL [2] utilised GAN to generate data rather than save raw to achieve rehearsal-free learning from class-incremental data. Despite progress in the class-incremental scenario of FCL, there remains a notable gap in the domain-incremental setting, which focuses on learning domain-invariant knowledge to reduce forgetting, similar to domain adaptation [6], a different approach from FCIL.
@@
To enhance local domain-invariant knowledge learning through the sharing of diverse domain information via global prompts in FDIL settings, we introduce RefFil, a rehearsal-free framework that leverages global prompts to share diverse domain information. As shown in Fig. 2, RefFil operates in task stages where participants engage in different domains. The process starts with a feature extractor generating a feature map from the input image, followed by our Client-wise Domain Adaptive Prompt (CDAP) generator. This generator encodes prompts using the feature map and a task ID. These encoded prompts, combined with the feature map, undergo classification involving computation of the cross-entropy loss and the Domain-specific Prompt Contrastive Learning (DPCL) loss with global prompts. Additionally, the feature map is merged with global prompts to compute the Global Prompt Learning (GPL) loss. All local prompts are sent to the central server to facilitate domain information sharing. For all preliminaries, including FDIL problem definition, please refer to Appendix A.
3 The Proposed Framework RefFil
Client-wise Domain Adaptive Prompt Generator. Most of the stages in FDIL training typically involve datasets from various domains, creating a substantial domain gap between participants’ local data. To navigate this challenge, generating refined prompts is essential for enabling participants to effectively distinguish and share prompts globally. Common approaches include maintaining a prompt pool, as seen in methods like L2P [37] and DualPrompt [36]. These strategies are similar to traditional continual learning methods, which rely on storing raw data or summarised representative information. However, a significant obstacle in FL is the limited memory resources of participant devices and privacy sensitivity. Additionally, as noted in the [16], a fixed-size prompt pool may not effectively encode domain-specific knowledge, and a scalable pool faces feasibility issues due to uncertainties in task requirements and computational constraints. This makes maintaining an extensive and adaptable prompt pool challenging for continuous learning in a federated context, where the unpredictability of tasks and resource limitations of devices must be considered.
Thus, inspired by the prompt encoder network structure proposed in [16], we propose a client-wise domain adaptive prompt generator. This generator is designed to create personalized prompts for each participant’s local data, embedding instance-level, fine-grained instructions with domain adaptation knowledge for global sharing. The CDAP generator $\mathcal{G}$ comprises several components: Layer Normalization (LN), a Multi-Layer Perceptron (MLP), a Cross-Client Domain Adaptation (CCDA) layer, and a Linear Transformation Layer (LT). This configuration enables the generation of adaptive local prompts $p_{m}$ for each participant $m$, which could be formulated as:
@@ 3 The Proposed Framework RefFil
where $\mathbf{I}$ is a set of input tokens, $\phi$ represents a linear layer that predicts two types of affine transformation parameters, $\alpha$ and $\lambda$, from a conditional input embedding $v$. Initially, $v$ is normalized, transposed, from $(n+1)\times d$ to $d\times(n+1)$ and then passed through an MLP to generate $p$-dimensional outputs which create current task domain well-fitted and instance-level prompts tailored with domain-specific knowledge. To handle the data that vary significantly in distribution from various domains, we introduce a CCDA layer, enhancing the CDAP generator’s generalization and robustness via FL aggregation. Additionally, we incorporate a Feature-Wise Linear Modulation (LT) [28] framework to augment the generated prompts with additional instructional guidance for prediction. LT achieves this by applying an affine transformation to the instance-level prompts, using scaling $\alpha_{v}$ and shifting parameters $\lambda_{v}$, both conditioned on the embedding $v$ as $[\alpha_{v},\lambda_{v}]=\phi(v)$. For the conditional input embedding $v$, we designed a task-specific key embedding layer that is also employed to encode local task IDs, linking tasks with domain-specific data.
@@
However, generally, a significant portion of participants, 80% in our scenario, transition to training tasks in new domains, while a minority continues with old tasks. This leads to a domain-specific prompt imbalance, with most prompts coming from new domains. This imbalance may bias local model updates towards these majority domains.
@@
where $c_{m}^{k}$ denotes the nearest neighbor (with the highest cosine similarity) of the class $k$ prompt from the $m$-th participant. Based on this clustering (Eq.4), we select representative prompts for each class from the embedding space. The final global clustered prompts are denoted as $\hat{P}^{g}$, are obtained as follows:
@@
Personalised Prompt Generator Optimization. We expect to improve the local model’s ability to differentiate between semantically related and unrelated prompts, ensuring a clear class decision boundary by enhancing sensitivity to domain-specific nuances in prompts. Specifically, we encourage the generation of prompts that are closely aligned with a semantically similar prompts set ($P^{+}_{i}$) while being distinctly different from a semantically unrelated prompts set ($P^{-}_{i}$). The prompts sampling strategy is defined as follows: For each class $k$, global prompts for class $k$ are selected, and their cosine similarity with the locally generated prompt of $k$ is computed. For clients $\mathcal{U}_{o}$ and $\mathcal{U}_{n}$ (single domain), the closest prompt is the positive sample $P^{+}$. For clients $\mathcal{U}_{b}$ (two domains), the two closest prompts are positive samples, and all remaining prompts are negative samples $P^{-}_{i}$. Drawing inspiration from the FPL [15], where contrastive learning was used to augment the model’s proficiency in learning domain-invariant knowledge, we introduce domain-specific prompt contrastive learning with temperature decay specifically designed to enhance the adaptability of prompt generation in the CDAP. It is natural to derive the following optimization objective term:
$\mathcal{L}_{DPCL}=-\log\frac{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})}{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})+\exp(\text{sim}(u_{i},P^{-}_{i})/\tau^{\prime})}\text{.}$ (6)
@@
Our Algorithm 1 of the FDIL process is outlined in Appendix B.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our framework on four widely used datasets for image classification tasks. The first, Digits-five [32, 33, 4], consists of 215,695 images across 10 classes and five domains, each image sized at 32x32 pixels. The second dataset, OfficeCaltech10 [35], contains 2,533 images distributed over 10 classes and four domains, also in 32x32 pixels. The third, PACS [19], includes 9,991 images across seven classes and four domains, each image being 224x224 pixels. Lastly, we use a subset of DomainNet [27], which we have named FedDomainNet, featuring 100,361 images in 48 classes across six domains, with a resolution of 224x224 pixels. All details of datasets can be found in Appendix C

### v2

@@ Abstract
Index Terms:
Federated Learning, Distributed Machine Learning, Continual Learning, Edge AI
I Introduction
@@
Federated Learning (FL) is a remarkable approach that enables various participants, also referred to as clients, to collaboratively train a global machine learning model while keeping their data decentralized, maintaining privacy and leveraging diverse data sources [18]. A significant limitation of current FL research lies in its primary focus on scenarios with static data distributions [14, 11], which remain unchanged throughout the training process [5, 4]. This limitation is particularly pronounced, as it fails to reflect the dynamic nature of real-world data. Recent studies have therefore shifted towards Federated Continual Learning (FCL), especially in class-incremental data scenarios [5, 4, 29, 17, 2], which struggles significantly with the critical problem of catastrophic forgetting [23], where model rapidly forgets previously learned knowledge upon encountering new data. In this paper, we explore more challenging domain-incremental data scenarios.
@@
Various methods have been proposed to address catastrophic forgetting. Rehearsal-based approaches, as described by Smith et al. [23], mitigate forgetting by replaying a selection of old task data during new task training. Alternatively, regularization-based methods, like Knowledge Distillation (KD) [15, 23, 5], guide the model’s optimization using historical data or models. Network expansion techniques [12, 7, 22] adapt the model structure to incorporate new features but their practicality in memory-constrained FL environments is limited. Recent research like L2P [28] and DualPrompt [27] explores prompt-based, rehearsal-free domain-incremental learning. However, these methods still do not fully address the challenge of using global domain information to enhance local model robustness in Federated Domain-Incremental Learning (FDIL).
@@
Strategies for mitigating catastrophic forgetting in domain-incremental tasks differ from those employed in class-incremental scenarios, where substantial differences between classes enable models to form clearer decision boundaries. In contrast, domain-incremental contexts involve more subtle inter-domain variations, which challenge a model’s ability to effectively separate domains. As a result, the focus shifts toward acquiring domain-invariant knowledge—an essential step in enhancing model robustness. This becomes particularly critical when models are expected to generalize across a broad spectrum of diverse domains. Despite its importance, research in this area remains limited, particularly regarding the global utilization of heterogeneous domain information from all clients throughout their learning progression. Such utilization is crucial for mitigating catastrophic forgetting by improving local models’ ability to learn domain-invariant representations in FDIL and, more broadly, within the FCL paradigm.
Figure 1: Key steps of RefFiL framework. Left panel (a) shows the common setting in existing FCL works, characterized by a cliff-style data transition. Right panel (a) shows our approach, where a subset of participants gradually transitions to new tasks. Panel (b) provides an overview example of a key step in our methodology: the 1st participant processes new domain data using global prompts from the 2nd to $m$-th participants and local prompts, enhancing robustness by aligning the model’s predictions across diverse domain prompts as inputs.
@@
II Preliminaries
Federated Domain-incremental Learning. In standard domain-incremental learning [19], a series of sequential tasks are denoted as $\mathcal{T}=\{\mathcal{T}^{t}\}_{t=1}^{T}$, with $T$ indicating the total number of tasks. Each task $\mathcal{T}^{t}$, the $t$-th in the sequence, consists of $N^{t}$ data-label pairs, i.e., $\mathcal{T}^{t}=\left\{x_{i}^{t},y_{i}^{t}\right\}_{i=1}^{N^{t}}$. Although each task is associated with a distinct domain $\mathcal{D}^{t}$, all tasks share the same label space $\mathcal{Y}$. The complete domain space is defined as $\mathcal{D}=\{\mathcal{D}^{t}\}_{t=1}^{T}$, with task domains differing in data distribution, i.e., $\mathcal{D}^{t}\neq\mathcal{D}^{t-1}$.
We adapt this setting to the FDIL scenario. Here, $M$ local participants, denoted as =$\mathcal{M}=\left\{\mathcal{M}_{m}\right\}_{m=1}^{M}$, interact with a central global server $\mathcal{M}_{G}$. During each global communication round $r$ (where $r=1,\dots,R$), a subset of local participants is randomly selected for gradient aggregation. When a participant $\mathcal{M}_{l}$ is selected for the $t$-th incremental task, it receives the latest global model $\theta^{{r},t}$.
The participant then trains this model on its local dataset $\mathbf{D}^{t}\sim\mathcal{P}_{l}^{\left|\mathbf{D}^{t}\right|}$, specific to task $t$, where $\mathcal{T}_{m}^{t}=\left\{x_{mi}^{t},y_{mi}^{t}\right\}_{i=1}^{N_{l}^{t}}\subset\mathcal{T}^{t}$ consists of data from new domains, and $\mathbf{P}_{m}$ denotes the data quantity distribution of participant $m$. These local datasets $\left\{\mathbf{P}_{m}\right\}_{m=1}^{M}$ are non-independent and identically distributed (non-iid), exhibiting a form of quantity shift. Each incremental task includes $K^{t}$ classes, maintaining consistent class counts ($K^{1}=K^{2}=\dots=K^{t}$) and labels ($\mathbf{y}^{1}=\mathbf{y}^{2}=\dots=\mathbf{y}^{t}$), but domain and data distributions vary across tasks ($\mathbf{x}^{1}\neq\mathbf{x}^{2}\neq\dots\neq\mathbf{x}^{t}$). Once $\mathcal{M}_{m}$ completes training on task $t$ with model $\theta^{r,t}$, it obtains an updated model $\theta^{r,t}_{l}$. All selected participants transmit their updates to the global server $\mathcal{M}_{G}$, where aggregation produces the updated global model $\theta^{(r+1),t}$. This model is then broadcast back to all participants, initiating the next communication round.
Client increment strategy. We divide participants into three dynamic groups for each incremental task: Old $\mathcal{U}_{o}$, In-between $\mathcal{U}_{b}$, and New $\mathcal{U}_{n}$. $\mathcal{U}_{o}$, with $M_{o}$ members, accesses only past domain data. $\mathcal{U}_{b}$ (with $M_{b}$ clients) works on both new and old domain data, while $\mathcal{U}_{n}$, consisting of $M_{n}$ new participants, focuses solely on new domain data from the current task. These groupings are redefined at each global round, with $\mathcal{U}_{n}$ being irregularly introduced during the FDIL process. As tasks progress, this strategy gradually increases the total number of participants, $M=M_{o}+M_{b}+M_{n}$.
Learning with Prompts. Prompt techniques, originally developed for task adaptation in natural language processing [16], have been primarily applied to large transformer models. However, they can also be adapted to convolutional neural network (CNN) based architectures, making them suitable for resource-constrained federated learning environments.
In our case, the model backbone includes a CNN-based feature extractor $h$, a patch embedding layer, attention block $b$, and a classifier $G$. Given an input image $x$, the feature map is obtained as $F=h(x)$. This is split into $n$ $d$-dimensional patch tokens $\mathrm{PT}$, and a trainable [CLS] token is prepended, forming the input token sequence:
This token sequence is then fed into the $b$-th attention block, producing:
where LN is layer normalization, $\mathbf{I}_{b}^{\prime}=\operatorname{LN}(\operatorname{MHSA}(\mathbf{I}_{b}))$ applies multi-head self-attention (MHSA), and $\mathbf{I}_{b}^{\prime\prime}=\operatorname{MLP}(\mathbf{I}_{b}^{\prime})$ is a multilayer perceptron (MLP). Each attention block consists of an MHSA layer, followed by MLP, skip connections [8], and LN [1] (see Fig. 2). The classifier $G$ maps the final [CLS] token to the class prediction:
$\hat{y}=G([\mathrm{CLS}]_{B}),$ (3)
where $\hat{y}$ denotes the predicted class probability distribution.
III Related Work
@@
Most recently, several works have begun to concentrate on FCL, especially Federated Class-incremental Learning (FCIL), which involves sequentially learning a series of distinct, non-overlapping classes. The primary challenge in FCIL is catastrophic forgetting, where models tend to lose previously learned knowledge upon task transition. Approaches such as GFLC [5] and its extension LGA [4] use knowledge distillation, employing an old model to regularize the current model’s output and a loss function to mitigate global forgetting. Method TARGET [29] uses a global model to generate synthetic data with a global distribution for the current training task to reduce forgetting. CFed [17] addresses data unavailability by generating pseudo-labels on a surrogate dataset for knowledge distillation and proposes a server distillation mechanism to mitigate intra-task forgetting. These methods typically require storing old tasks or extra data, which is impractical in FL with resource-limited devices. MFCL [2] utilised GAN to generate data rather than save raw to achieve rehearsal-free learning from class-incremental data. Despite progress in the class-incremental scenario of FCL, there remains a notable gap in the domain-incremental setting, which focuses on learning domain-invariant knowledge to reduce forgetting, a different approach from FCIL.
IV The Proposed Framework RefFiL
@@
To enhance local domain-invariant knowledge learning through the sharing of diverse domain information via global prompts in FDIL settings, we introduce RefFiL, a rehearsal-free framework that leverages global prompts to share diverse domain information. As shown in Fig. 2, RefFiL operates in task stages where participants engage in different domains. The process starts with a feature extractor generating a feature map from the input image, followed by our Client-wise Domain Adaptive Prompt (CDAP) generator. This generator encodes prompts using the feature map with the task ID. These encoded prompts, combined with the feature map, undergo classification involving computation of the cross-entropy loss and the Domain-specific Prompt Contrastive Learning (DPCL) loss with global prompts. Additionally, the feature map is merged with global prompts to compute the Global Prompt Learning (GPL) loss. All local prompts are sent to the central server to facilitate domain information sharing.
Client-wise Domain Adaptive Prompt Generator. Most of the stages in FDIL training typically involve datasets from various domains, creating a substantial domain gap between participants’ local data. To navigate this challenge, generating refined prompts is essential for enabling participants to effectively distinguish and share prompts globally. Common approaches include maintaining a prompt pool, as seen in methods like L2P [28] and DualPrompt [27]. These strategies are similar to traditional continual learning methods, which rely on storing raw data or summarised representative information. However, a significant obstacle in FL is the limited memory resources of participant devices and privacy sensitivity. Additionally, as noted in [10], a fixed-size prompt pool may not effectively encode domain-specific knowledge, and a scalable pool faces feasibility issues due to uncertainties in task requirements and computational constraints. This makes maintaining an extensive and adaptable prompt pool challenging for continuous learning in a federated context, where the unpredictability of tasks and resource limitations of devices must be considered.
Thus, inspired by prompt encoder network structure proposed in [10], we propose a client-wise domain adaptive prompt generator. This generator is designed to create personalized prompts for each participant’s local data, embedding instance-level, fine-grained instructions with domain adaptation knowledge for global sharing. The CDAP generator $\mathcal{G}$ comprises several components: Layer Normalization (LN), a Multi-Layer Perceptron (MLP), a Cross-Client Domain Adaptation (CCDA) layer, and a Linear Transformation Layer (LT). This configuration enables the generation of adaptive local prompts $p_{m}$ for each participant $m$, which could be formulated as:
@@
where $\mathbf{I}$ is a set of input tokens, $\phi$ represents a linear layer that predicts two types of affine transformation parameters, $\alpha$ and $\lambda$, from a conditional input embedding $v$. Initially, $v$ is normalized, transposed, from $(n+1)\times d$ to $d\times(n+1)$ and then passed through an MLP to generate $p$-dimensional outputs which create current task domain well-fitted and instance-level prompts tailored with domain-specific knowledge. To handle the data that vary significantly in distribution from various domains, we introduce CCDA layer, a globally transferable linear layer enhancing the CDAP generator’s generalization and robustness via FL aggregation. Additionally, we incorporate a Feature-Wise Linear Modulation (LT) [21] framework to augment the generated prompts with additional instructional guidance for prediction. LT achieves this by applying an affine transformation to the instance-level prompts, using scaling $\alpha_{v}$ and shifting parameters $\lambda_{v}$, both conditioned on the embedding $v$ as $[\alpha_{v},\lambda_{v}]=\phi(v)$. For the conditional input embedding $v$, we designed a task-specific key embedding layer that is also employed to encode local task IDs, linking tasks with domain-specific data. However, the task ID contributes only to training process by helping to distinguish data from different tasks, and is not utilized during the inference stage.
@@
However, generally, a significant portion of participants, 80% in our scenario, transition to training tasks in new domains, while a minority continues with old tasks. This leads to a domain-specific prompt imbalance, with most prompts from new domains. This imbalance may bias local model updates towards these majority domains.
@@
where $c_{m}^{k}$ denotes the nearest neighbor (with the highest cosine similarity) of the class $k$ prompt from the $m$-th participant. Based on this clustering (Eq.7), we select representative prompts for each class from the embedding space. The final global clustered prompts are denoted as $\hat{P}^{g}$, are obtained as:
@@
Personalised Prompt Generator Optimization. We expect to improve the local model’s ability to differentiate between semantically related and unrelated prompts, ensuring a clear class decision boundary by enhancing sensitivity to domain-specific nuances in prompts. Specifically, we encourage the generation of prompts that are closely aligned with a semantically similar prompts set ($P^{+}_{i}$) while being distinctly different from a semantically unrelated prompts set ($P^{-}_{i}$). The prompts sampling strategy is defined as follows: For each class $k$, global prompts for class $k$ are selected, and their cosine similarity with the locally generated prompt of $k$ is computed. For clients $\mathcal{U}_{o}$ and $\mathcal{U}_{n}$ (single domain), the closest prompt is the positive sample $P^{+}$. For clients $\mathcal{U}_{b}$ (two domains), the two closest prompts are positive samples, and all remaining prompts are negative samples $P^{-}_{i}$. Drawing inspiration from FPL [9], where contrastive learning was used to augment the model’s proficiency in learning domain-invariant knowledge, we introduce domain-specific prompt contrastive learning with temperature decay specifically designed to enhance the adaptability of prompt generation in the CDAP. It is natural to derive following optimization objective term:
$\mathcal{L}_{DPCL}=-\log\frac{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})}{\exp(\text{sim}(u_{i},P^{+}_{i})/\tau^{\prime})+\exp(\text{sim}(u_{i},P^{-}_{i})/\tau^{\prime})}\text{.}$
@@
Our FDIL process is outlined in Algorithm 1.
Output : Trained global model $\theta^{R}$
1
2Central Server
3       Initialise global model $\theta^{r}$
4      for $t=1$ to $T$ do
5             for $r=1$ to $R$ do
6                   Participants Random Selection
7                  Broadcast $\theta^{r,t},\hat{P}^{g}$rightarrow Selected Participants
8                  $\theta^{r}_{m}$, $LPG^{k}_{m}\leftarrow$ Selected Participants
11                  $\hat{P}^{g}=\left[\hat{P}^{g,1},\dots,\hat{P}^{g,k}\right]$
12
14
15
16Participants
18      $\theta^{r,t},\hat{P}^{g}$leftarrow Central Server
[…]
20      if $m$ in $\mathcal{U}_{b}$ then
21             $\mathbf{D}^{t}_{m}=\text{concat}(\mathbf{D}^{t-1}_{m},\mathbf{D}^{t}_{m})$
22
23      for $e=0,...,E$ do
24             for $x_{i},y_{i}\in\mathbf{D}^{t}_{m}$ do
25                   $u_{i},p_{l},\xi_{l},\xi_{g}=\theta^{r,t}(x_{i},\bar{P}^{g})$
26                  $\mathcal{L}_{CE}\leftarrow$ Eq.13 $(u_{i},\xi_{l})$
27                  $\mathcal{L}_{GPL}\leftarrow$ Eq.12 $(u_{i},\xi_{g})$
28                  $\mathcal{L}_{DPCL}\leftarrow$ Eq.9 $(p_{l},\hat{P}^{g})$
29                  $\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{GPL}+\mathcal{L}_{DPCL}$
31                  if $e=(E-1)$ then
32                         $P_{m}[k]$ append $p_{i}$
33
34
35
36      for $k$ in $P_{m}$ keys do
38
39      Send $\theta^{r}_{m}$, $LPG_{m}$rightarrow Central Server
Algorithm 1 Rehearsal-free federated domain-incremental learning
V Experiments
V-A Experimental Setup
Datasets. We evaluate our framework on four widely used datasets for image classification tasks. The first, Digits-five [25], consists of 215,695 images across 10 classes and five domains, each image sized at 32x32 pixels. The second dataset, OfficeCaltech10 [26], contains 2,533 images distributed over 10 classes and four domains, also in 32x32 pixels. The third, PACS [13], includes 9,991 images across seven classes and four domains, each image being 224x224 pixels. Lastly, we use a subset of DomainNet [20], which we have named FedDomainNet, featuring 100,361 images in 48 classes across six domains, with a resolution of 224x224 pixels.